OpenStudy (anonymous):

[6.02] Fun Land theme park has a total of 80 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 80 60 30 20 can any one can help me

OpenStudy (anonymous):

Is it 80

jimthompson5910 (jim_thompson5910):

let m = # of males and f = # of females

jimthompson5910 (jim_thompson5910):

Fun Land theme park has a total of 80 males and females which means m + f = 80 solve for f to get f = 80 - m

jimthompson5910 (jim_thompson5910):

If the number of males is 20 more than twice the number of females tells us that m = 2f + 20

jimthompson5910 (jim_thompson5910):

now plug in f = 80 - m to get m = 2f + 20 m = 2(80 - m) + 20 m = 160 - 2m + 20

jimthompson5910 (jim_thompson5910):

I'll let you finish up

OpenStudy (anonymous):

i will =138 loli don't know how to solve it :( thanks for helping me

jimthompson5910 (jim_thompson5910):

m = 160 - 2m + 20 m = 180 - 2m m+2m = 180 3m = 180 m = 180/3 m = 60

OpenStudy (anonymous):

i see now

OpenStudy (anonymous):

can you help on this. [6.03] What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

jimthompson5910 (jim_thompson5910):

double everything in equation 2 to get -4x+6y = 12 now add the equations to get 3x - 6y = 12 -4x+6y = 12 ------------- -x + 0y = 24 -x = 24 so x = ??? Note: this is NOT the answer since you need to find y, not x

OpenStudy (anonymous):

Do i x times 24 how do i solve

jimthompson5910 (jim_thompson5910):

if -x = 24, then x = -24

jimthompson5910 (jim_thompson5910):

now use this to find y

OpenStudy (anonymous):

x=1

jimthompson5910 (jim_thompson5910):

-2x + 3y = 6 -2(-24) + 3y = 6 48 + 3y = 6 solve for y

OpenStudy (anonymous):

so 48 + 3y =6 -3 -3 48 = 3 48 48 right

jimthompson5910 (jim_thompson5910):

48 + 3y = 6 48 + 3y-48 = 6-48 3y = -42 y = ???

OpenStudy (anonymous):

Y= -14

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

want to help on this two [6.06] Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -2x + 2 (-4, 0) (-3, 2) (2, -1) (0, 6)

jimthompson5910 (jim_thompson5910):

plug each point into both inequalities and see which point makes them both true

jimthompson5910 (jim_thompson5910):

the point that makes both true is a solution

OpenStudy (anonymous):

so how do i sit it up

jimthompson5910 (jim_thompson5910):

for example y > x + 4 0 > -4 + 4 .... plug in (-4,0) 0 > 0... that's false, so choice A is out

OpenStudy (anonymous):

can you give me a example so we can both solve?

jimthompson5910 (jim_thompson5910):

try out choice B and show me what you get

OpenStudy (anonymous):

(-3, 2) y<-3+2 i need help now

jimthompson5910 (jim_thompson5910):

y > x + 4 y > -3 + 4 ... plug in x = -3 2 > -3 + 4 ... plug in y = 2 keep going

OpenStudy (anonymous):

how you got 4 did i do something wrong??

jimthompson5910 (jim_thompson5910):

the 4 is part of the first inequality

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I am lost now

jimthompson5910 (jim_thompson5910):

you basically start with y > x+4 then plug in (-3,2)

OpenStudy (anonymous):

ok so what is next

jimthompson5910 (jim_thompson5910):

simplify

OpenStudy (anonymous):

so simplify the y>x+4 then plug in -3 2

jimthompson5910 (jim_thompson5910):

no, plug it in first, then simplify

OpenStudy (anonymous):

oh ok so it will be like this -3y+4 right???

jimthompson5910 (jim_thompson5910):

you just simplify this 2 > -3 + 4

OpenStudy (anonymous):

oh ok sorry

OpenStudy (anonymous):

so = 9?

jimthompson5910 (jim_thompson5910):

2 > -3 + 4 2 > 1 is that true? or false?

OpenStudy (anonymous):

False

jimthompson5910 (jim_thompson5910):

2 > 1 is true since 2 is indeed larger than 1

OpenStudy (anonymous):

oh